if x is N(0,4) Find P(1<=x<=2) and P(1<=x<=2 | x>=1)
Answer by stanbon(75887)

if x is N(0,4)
Find P(1<=x<=2)
---
z(1) = (1-0)/4 = 1/4
z(2) = (2-0)/4 = 1/2
---
P(1<= x <=2) = P(1/4 < z < 1/2) = 0.0928
------

and P(1<=x<=2 | x>=1)
= P(1<=x<=2 AND x>=1)/P(x>=1)
= P(1<= x <=2)/P(x>= 1)
---
P(x>=1) = P(z >= 1/4) = 0.4013
---
Ans: 0.0928/0.4013 = 0.2313
===========================
Cheers,
Stan H.
==========